OpenStudy (anonymous):

What is the equation of the line passing through (2, –2) and perpendicular to the line y = –x + 3 in slope-intercept form?

OpenStudy (anonymous):

y = -x . Same slope, plug in x = 2, you get -2 = y.

OpenStudy (anonymous):

\[y-yo=m(x-xo)\] Using it on the line you gave \[y-0=-1(x-3)\] so you know the m (slope) is equal to -1 Replacing to the point you want: \[y+2=-1(x-2)\] \[y+2=-x+2\] \[y+x=0\]

OpenStudy (phi):

You should memorize this rule. two lines are perpendicular if slope of line1 = m and the slope of line2= -1/m Also, they are parallel if they have the same slope m. In your case, m= -1, so the slope of the 2nd line is -1/-1 = 1

OpenStudy (phi):

Once you know the slope, use Luis... formula (y -y0) = m(x-x0) where (x0,y0) is any point on your line. In your case (2,-2)

OpenStudy (phi):

Finally, put it into slope-intercept form y= mx+b